# Do graph neural network states contain graph properties?

### v1

@@ 7.4 Future work
Our methodology has several limitations. While we addressed dataset issues such as leakage and isomorphic graphs, a key challenge remains the lack of guarantees that GNNs find globally optimal solutions, despite their theoretical capacity as universal function approximators. This is particularly evident in fMRI data, where multiple layers of complexity—from MRI limitations and BOLD signal characteristics to Pearson correlation for functional connectivity—introduce noise and inaccuracies. Investigating additional graph properties like girth or complex motifs could be beneficial. Preliminary work on alternative architectures (e.g., GATv2, GraphSAGE, ChebNet, Set2Set, HO-Conv, DiffPool) has begun but is not yet complete.
@@ 8 Conclusion
We demonstrate the relevance of our model-agnostic explainability method for graph neural networks which probe for GNN graph theoretic representations on the Grid-House dataset. We anticipate any lazy learning bias. We manifest both the expressivity of different GNN architectures and their ability to solve a graph classification problem through optimal feature extraction. They render it linearly separable in the space of their embeddings through the computation of the number of squares in the graph.
That experiment prompted us to investigate both the Clintox Molecular dataset and fMRI FC connectomes dataset and anchored the possibility of formulating hypotheses on the emergent dependence of complex systems qualities to basic and more higher level structural properties. This kind of higher order properties are very interesting in order to understand how the macroscopic behavior of complex systems emerges from the intricate interplay of their microscopic components. For example how diseases spread in social networks, how neurons interact in the brain, or how information propagates through the Internet. There is a manifest emergence of molecular qualities like toxicity with regard to their structural properties like node degree (atom valency) and spectral radius (the molecule’s stability). But the complexity of complex systems like the brain makes blurrier the possibility of understanding what affects what as, for example, one could argue that behavioural therapies might influence the brain connectivity as the brain connectivity might influence behavioural qualities of one patient. Echoing this egg-chicken conundrum, the investigation of motifs and spectral properties’ role in neurological disorders like ASD and MDD could allow for promising avenues.
@@
Pelletreau-Duris Tom: Conceptualization (lead); methodology (lead); software (lead); formal analysis (lead); writing – original draft (lead); writing – review and editing (equal). Ruud van Bakel: writing – review and editing (equal); Supervision Michael Cochez:Writing – review and editing (equal); Supervision; Funding acquisition
Acknowledgments
RvB and MC are in part funded by the Graph-Massivizer project, the Horizon Europe research and innovation program of the European Union grant management number 101093202 https://graph-massivizer.eu/
@@ Appendix A R2 Score
From an information-theoretic perspective, training the probing classifier $g$ can be viewed as estimating the mutual information between the learned representations $f_{l}(x)$ and the property $z$. This mutual information is denoted as $\mathrm{I}(\mathbf{z};\mathbf{h})$, where $\mathbf{z}$ refers to the property and $\mathbf{h}$ represents the intermediate representations Belinkov (2021).
@@ Appendix B Local and global graph properties
B.1 Artificial dataset
B.1.1 Grid House figures
@@
B.1.2 Grid House models
@@
B.1.3 Grid House RESULTS
B.1.4 Graph properties probing results
@@
B.1.5 Graph properties probing plots
@@
Figure 5: T-SNE visualisation across different layers of our GCN architecture aligned with the probing $R^{2}$ scores plots (Grid House)
Figure 6: Plot of the GCN (control) $R^{2}$ results across different layers probing for graph properties with post pooling layers only (Grid House)
Figure 7: Plot of the GCN ($L_{2}$) $R^{2}$ results across different layers probing for graph properties with post pooling layers only(Grid House)
Figure 8: Plot of the GCN (dropout) $R^{2}$ results across different layers probing for graph properties with post pooling layers only(Grid House)
@@
Figure 9: Plot of the GIN (control) $R^{2}$ results across different layers probing for graph properties with post pooling layers only (Grid House)
Figure 10: Plot of the GIN ($L_{2}$) $R^{2}$ results across different layers probing for graph properties with post pooling layers only (Grid House)
Figure 11: Plot of the GIN (dropout) $R^{2}$ results across different layers probing for graph properties with post pooling layers only (Grid House)
@@
Figure 12: Plot of the GAT $R^{2}$ results across different layers probing for graph properties with post pooling layers only (Grid House)
B.1.6 Grid House Node properties probing results
@@
B.2 Clintox dataset
B.2.1 model
@@ B.2 Clintox dataset
B.2.2 Results
B.2.3 Graphs properties probing results
@@ B.2.4 Plots
B.2.5 Node properties probing results
@@
B.3 fMRI datasets
B.3.1 models
@@
B.3.2 RESULTS ABIDE (ASD) dataset
@@
B.3.3 RESULTS REST-meta-MDD (MDD) dataset
@@
B.3.5 Results ASD and MDD node properties
B.3.6 Node properties probing results
@@
Appendix C Brain Imaging and GNNs
Our brain is a network, more precisely a complex network of functionally interconnected regions specialised in specific cognitive tasks, sharing information with each other. In the last three decades, the field of biological neuroscience and computational cognitive neuroscience have provided and incredible amount of knowledge on the role, function and biological structure of such regions of interests, aiming at better understanding both the biological organisation of the brain (which we can refer to as the ’hardware implementation’), the representation embedded in this hardware and the computational strategy employed to treat this kind of representation Marr (1984). In other terms, we got better at understanding how each region independently organises itself and processes and forms information (cite Connecting network science and information theory). The main problem for modern computational neuroscience consists of understanding the brain’s plasticity (how regions change over time), the inter-individual differences (how regions specialise differently between people) and how the brain integrates the information (how regions communicate with regard to each other).
For example we understand very well more basic brain structures like the cerebellum due to its high inter-individual similarity but we have a lot more difficulties modelling the prefrontal cortex which is so different from an individual to the other Kanai & Rees (2011); Gu & Kanai (2014); Mills et al. (2021). In other terms, we do understand well the brain operating in segregation but not so much in integration Aine (1995). Functional segregation refers to the distinct specialisation of anatomical brain regions and functional integration refers to the possible temporal dependencies between the activity of anatomically separated regions of the brain.
@@ Appendix C Brain Imaging and GNNs
Then, relying on fMRI, we have several ways to study the functional connectivity of the brain. Functional connectivity is defined as the temporal dependence of neuronal activity patterns of anatomically separated brain regions Aertsen et al. (1989); Friston et al. (1993) and studies have shown that we could study functional connectivity between brain regions as the level of coactivation of functional MRI time-series Lowe et al. (1998; 2000). As a result, conceptualising the brain as an integrative network of functionally interacting brain regions offers a powerful framework for understanding large-scale neuronal communication. It provides a method for investigating how functional connectivity and information integration relates to human behaviour and how this organisation may be altered in neurodegenerative diseases Bullmore & Sporns (2009); Greicius et al. (2009).
@@
To understand how a specific brain region interacts with others, researchers most often analyse its resting-state activity and use simple pearson correlation of time-series data of a region with the time-series data of all other brain regions, they create a functional connectivity map (fcMap), which visually represents the strength of these connections Biswal et al. (1997); Cordes et al. (2000). This is basically a matrix with value and we can understand it as a non relational data structure, in other terms, a graph.
@@
More and more work in cognitive neurosciences explore the link between graph theory and connectomes (functional connectivity matrices) Farahani et al. (2019). By representing brain regions as nodes and their connections as edges, graph theory provides a powerful framework for analysing the structural and functional organisation of the brain. Notably, studies have begun to explore the link between structural properties of brain connectivity, as captured by connectomes, and the manifestation of neurological disorders such as Autism Spectrum Disorders (ASD) and Major Depressive Disorders (MDD). ASD, characterised by impairments in social communication and repetitive behaviours. MDD is characterised as a mood disorder marked by persistent sadness and loss of interest.
@@
These findings highlight the potential of connectome analysis to elucidate the neurological underpinnings of NDs and pave the way for the development of novel diagnostic and therapeutic strategies. Studying the link between the brain’s Functional connectivity signature and behavioural quality of patients through probing learned embeddings of neural networks trained on classification tasks could thus be a promising avenue to help disentangle the gap between its segregational characteristics and the emergence Johnson (2002); Eccles (1994); Wang et al. (2015); Carroll & Parola (2024) of higher level behavioural quality.
@@
However, if NDs result in alterations in brain functional and structural connections, as well as local and global connections Seeley et al. (2009); Wang et al. (2015); Pasquini et al. (2015); Stam et al. (2007), traditional deep learning models such as CNN and LSTM are difficult to fit to the connectivity of the brain Zhang et al. (2023). These long range dependencies, though, are well captured by the relational models defined previously in this thesis : Graph Neural Networks.
@@
Definition : Psychiatric diagnosis can be regarded as a graph classification task. Given an input graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with node feature matrix $X$, GNNs employ the message-passing paradigm to propagate and aggregate the representations of information along edges to generate a node representation $h_{v}$ for each node $v\in\mathcal{V}$ and then explore the modelled human brains using graph methods to extract abnormal brain networks, subnetworks, and local connections Palop et al. (2006); Thomas et al. (2016).
@@
Similarly to Zheng et al. (2023), a GNN can be formally defined through an aggregation function A and a combine function C such that $h_{v}^{(k)}$ is the node embedding of node $v$ at the $k$-th layer and $\mathcal{N}(v)$ is the set of neighbour nodes of $v$:
@@
In the context of connectomes, many studies have focused on the relationship between general intellectual ability and small-world characteristics in intrinsic functional networks for describing individual differences in general intelligence van den Heuvel & Hulshoff Pol (2010); van den Heuvel et al. (2009); Langer et al. (2012); Hilger et al. (2017). Better intellectual performance was associated with shorter characteristic path length, the nodal centrality of hub regions in the salience network, as well as the efficiency of functional integration between the frontal and parietal areas Jung & Haier (2007) In general, when connections between specialised brain regions are disrupted, even within localised areas, the result is often functional impairment. This impairment is linked to atypical integration of activity across distributed brain networks Ffytche & Catani (2005); Catani et al. (2005). Characterising this impairment through the use of GNN could be one application of our probing pipeline. So far, GNNs have achieved promising diagnostic accuracy on autism spectrum disorder (ASD) Rakhimberdina et al. (2020), schizophrenia Rakhimberdina & Murata (2020), bipolar disorder (BD) Yang et al. (2019) and MDD Zheng et al. (2023). We’ll focus on ASD and MDD. But here as in other graph related fields, research has highlighted we were lacking Interpretability Zheng et al. (2023).
For ASD, The contribution of rs-fMRI studies based on graph theory for autism exploration is important Redcay et al. (2013); Rudie et al. (2013); Di Martino et al. (2014); Keown et al. (2017); Kazeminejad & Sotero (2019). Studies have found increased short-range connections in ASD, particularly within sensory and association cortices. This local overconnectivity may contribute to the sensory sensitivities and restricted interests often seen in ASD. Conversely, long-range connections between distant brain regions tend to be reduced in ASD. This underconnectivity affects integration of information across brain networks. Based on this literature Farahani et al. (2019) we know that the modularity, clustering coefficient, and local efficiency are relatively reduced in ASD (i.e., inefficiency of information transmission in a particular module) while global communication efficiency is increased (shorter average path lengths). As another example, Redcay et al. (2013) observed an increase in betweenness centrality and local connections by analysing the prefrontal brain areas in adolescents with ASD.
@@
For patients with MDD, several studies have reported topological changes in human brain connectome, including a loss of the small-world network Ye et al. (2015); Achard & Bullmore (2007)] and a significant reorganisation of the community structure Zhang et al. (2011); Leistedt et al. (2009); Lord et al. (2012). In general, MDD patients exhibit increased global and local clustering coefficients, indicating a higher degree of local interconnectedness and efficiency in information processing. Moreover, increased modularity in MDD patients indicated that there were relatively less inter-modular edges and more intra-modular edges, which may also be associated with the disruptions in emotion regulation by decreasing communications between the Default Mode Network (DMN) and the Cognitive Control Network (CCN) Ye et al. (2015). We would thus expect that classifying FC matrices with regard to MDD should use more clustering coefficient, clusterization properties and modularity measures than random (like the presence of motifs like the number small clusters, squares or triangles).

### v2

@@ 6.3 Future work
Our methodology has several limitations. While we addressed dataset issues such as leakage and isomorphic graphs, a key challenge remains the lack of guarantees that GNNs find globally optimal solutions, despite their theoretical capacity as universal function approximators. This is particularly evident in fMRI data, where multiple layers of complexity—from MRI limitations and BOLD signal characteristics to Pearson correlation for functional connectivity—introduce noise and inaccuracies. Investigating additional graph properties like girth or complex motifs could be beneficial. Preliminary work on alternative architectures (e.g., GATv2, GraphSAGE, ChebNet, Set2Set, HO-Conv, DiffPool) has begun but is not yet complete. An extensive exploration of 1-WL, 2-WL and 3-WL GNN equivalent could bolster the paper’s contributions by showing clear restrictions and capabilities of these models. As a future encouraging work, the question of supervision signal would be very insightful in understanding what type of properties are learn by unsupervised models.
@@ 7 Conclusion
We demonstrate the relevance of using probing classifier as a model-agnostic explainability method for graph neural networks. We manifest both the expressivity of different GNN architectures and their ability to solve a graph classification problem through optimal feature extraction. They render it linearly separable in the space of their embeddings through the computation of graph properties. We validate domain knownledge with the Clintox Molecular dataset and investigate the possibility of formulating hypotheses on the emergent dependence of complex systems attributes to basic and more higher level structural properties with the fMRI FC connectomes dataset. We explore how the macroscopic behavior of complex systems emerges from the intricate interplay of their microscopic components. There is a manifest emergence of molecular qualities like toxicity with regard to their structural properties like node degree (atom valency) and spectral radius (the molecule’s stability). To explain a macro attribute, there are instances where structural properties may offer more insight than the mere aggregation of element properties. Both can be interconnected, yet the challenge remains in disentangling these relationships.
@@
Pelletreau-Duris Tom: Conceptualization (lead); methodology (lead); software (lead); formal analysis (lead); writing – original draft (lead); writing – review and editing (equal). Ruud van Bakel: writing – review and editing (equal). Michael Cochez:Writing – review and editing (equal).
Code
All results and experiments accessible on github
@@ Appendix A R2 Score
$R^{2}=1-\frac{\sum_{i}\left(z^{(i)}-\hat{z}^{(i)}\right)^{2}}{\sum_{i}\left(z^{(i)}-\bar{z}\right)^{2}}$
Where: $z^{(i)}$ is the ground truth value of the property $\hat{z}$ for the $i$-th data point in the probing dataset $\mathcal{D}_{P}$. $\hat{z}^{(i)}$ is the predicted value of $\hat{z}$ produced by the probing classifier $g$. $\bar{z}$ is the mean of the ground truth values $z^{(i)}$ over the dataset. The numerator represents the residual sum of squares (how far off the predictions are), and the denominator represents the total sum of squares (the variance in the ground truth values).
[…]
From an information-theoretic perspective, training the probing classifier $g$ can be viewed as estimating the mutual information between the learned representations $f_{l}(x)$ and the property $z$. This mutual information is denoted as $\mathrm{I}(\mathbf{z};\mathbf{h})$, where $\mathbf{z}$ refers to the property and $\mathbf{h}$ represents the intermediate representations (Belinkov, 2021).
@@ Appendix B Local and global graph properties
We are using the Small-World Index, $SWI=\left(\frac{L-L_{l}}{L_{r}-L_{l}}\right)\times\left(\frac{C-C_{r}}{C_{l}-C_{r}}\right)$ in our experiment because it provides a more balanced and robust measure of small-world properties. Unlike the Small-World Quotient: $Q=\frac{C/C_{r}}{L/L_{r}}$, which can be sensitive to network size and degree, $SWI$ normalizes both the clustering coefficient and average path length with respect to both random and lattice reference graphs. This dual normalization approach ensures that $SWI$ is less prone to false positives or negatives, making it a more reliable metric for our analysis (Neal, 2017).
Appendix C Literature review on related Work
Existing post-hoc GNN explanations methods can be classified into two main categories: instance-level and model-level methods (Barredo Arrieta et al., 2020). See (Agarwal et al., 2023; Dai et al., 2022) for nice reviews on the subject. In the realm of instance based methods, gradient-based methods use the gradients of the output with respect to the input or intermediate features to measure the importance of each component of the graph. Decomposition-based methods try to decompose the input graph into smaller subgraphs or paths that can account for the output. Surrogate-based methods use a simpler, more interpretable model to approximate the behavior of the original GNN and provide explanations based on the surrogate model. And finally Perturbation-based methods which perturb the input graph by removing or adding nodes, edges, or features, and observe the changes in the output to identify the influential components. The most mainstream technique, GNNExplainer (Ying et al., 2019) achieves explanation by removing redundant edges from an input graph instance, maximizing the mutual information between the distribution of subgraphs and the GNN’s prediction. It is able to provide an explanation both in terms of a subgraph of the input instance to explain, and a feature mask indicating the subset of input node features which is most responsible for the GNN’s prediction.
For model-based techniques, few methods come to mind (Saha et al., 2022; Azzolin et al., 2023; Vu & Thai, 2020; Wang et al., 2023; Xuanyuan et al., 2023; Yuan et al., 2020; Zhang et al., 2021). The most mainstream method seems to be XGNN (Yuan et al., 2020). The authors of XGNN investigate the possible input characteristics used by a GNN for graph classification. But they formulate the problem as a reinforcement learning problem and generate graph patterns iteratively. Such an iterative approach is often intractable for large graphs. Moreover, it does not allow for both node classification and graph classification explanations, nor does it allow for an investigation of the learning process through the different layers of the GNN. In general, none of the techniques allow for an interpretation of the hidden representations states with graph properties.
Appendix D Grid House artificial dataset
D.1 Grid House figures
@@
D.2 Grid House models
@@
Figure 6: Plot of the correlation between the different model test accuracies and their maximum R2 score (Grid House)
D.3 Grid House RESULTS
D.3.1 Graph properties probing results
@@
D.3.2 Graph properties probing plots
@@
Figure 7: T-SNE visualization across different layers of our GCN architecture aligned with the probing $R^{2}$ scores plots (Grid House)
Figure 8: T-SNE visualization across different layers of our GCN architecture aligned with the probing $R^{2}$ scores plots with mean-pooled node embeddings (Grid House)
Figure 9: Plot of the GCN (control) $R^{2}$ results across different layers probing for graph properties with post pooling layers only, allowing clearer visualization and higher order property interpretation (Grid House)
Figure 10: Plot of the GCN ($L_{2}$) $R^{2}$ results across different layers probing for graph properties with post pooling layers only, allowing clearer visualization and higher order property interpretation (Grid House)
Figure 11: Plot of the GCN (dropout) $R^{2}$ results across different layers probing for graph properties with post pooling layers only, allowing clearer visualization and higher order property interpretation (Grid House)
@@
Figure 12: T-SNE visualization across different layers of our GIN architecture aligned with the probing $R^{2}$ scores plots with mean-pooled node embeddings (Grid House)
Figure 13: Plot of the GIN (control) $R^{2}$ results across different layers probing for graph properties with post pooling layers only, allowing clearer visualization and higher order property interpretation (Grid House)
Figure 14: Plot of the GIN ($L_{2}$) $R^{2}$ results across different layers probing for graph properties with post pooling layers only, allowing clearer visualization and higher order property interpretation (Grid House)
Figure 15: Plot of the GIN (dropout) $R^{2}$ results across different layers probing for graph properties with post pooling layers only, allowing clearer visualization and higher order property interpretation (Grid House)
@@
Figure 16: T-SNE visualization across different layers of our GAT architecture aligned with the probing $R^{2}$ scores plots (Grid House)
Figure 17: T-SNE visualization across different layers of our GAT architecture aligned with the probing $R^{2}$ scores plots with mean-pooled node embeddings (Grid House)
Figure 18: Plot of the GAT $R^{2}$ results across different layers probing for graph properties with post pooling layers only, allowing clearer visualization and higher order property interpretation (Grid House)
D.3.3 Grid House Node properties probing results
@@
Appendix E Clintox dataset
E.1 model
@@ Appendix E Clintox dataset
E.2 Results
E.2.1 Graphs properties probing results
@@ E.2.2 Plots
E.2.3 Node properties probing results
@@
Appendix F fMRI datasets
F.1 models
@@
F.2 RESULTS ABIDE (ASD) dataset
@@
F.3 RESULTS REST-meta-MDD (MDD) dataset
@@
F.5 Results ASD and MDD node properties
@@
Appendix G Brain Imaging and GNNs
Our brain is a network, more precisely a complex network of functionally interconnected regions specialised in specific cognitive tasks, sharing information with each other. In the last three decades, the field of biological neuroscience and computational cognitive neuroscience have provided and incredible amount of knowledge on the role, function and biological structure of such regions of interests, aiming at better understanding both the biological organisation of the brain (which we can refer to as the ’hardware implementation’), the representation embedded in this hardware and the computational strategy employed to treat this kind of representation (Marr, 1984). In other terms, we got better at understanding how each region independently organises itself and processes and forms information (cite Connecting network science and information theory). The main problem for modern computational neuroscience consists of understanding the brain’s plasticity (how regions change over time), the inter-individual differences (how regions specialise differently between people) and how the brain integrates the information (how regions communicate with regard to each other).
For example we understand very well more basic brain structures like the cerebellum due to its high inter-individual similarity but we have a lot more difficulties modelling the prefrontal cortex which is so different from an individual to the other (Kanai & Rees, 2011; Gu & Kanai, 2014; Mills et al., 2021). In other terms, we do understand well the brain operating in segregation but not so much in integration (Aine, 1995). Functional segregation refers to the distinct specialisation of anatomical brain regions and functional integration refers to the possible temporal dependencies between the activity of anatomically separated regions of the brain.
@@ Appendix G Brain Imaging and GNNs
Then, relying on fMRI, we have several ways to study the functional connectivity of the brain. Functional connectivity is defined as the temporal dependence of neuronal activity patterns of anatomically separated brain regions (Aertsen et al., 1989; Friston et al., 1993) and studies have shown that we could study functional connectivity between brain regions as the level of coactivation of functional MRI time-series (Lowe et al., 1998; 2000). As a result, conceptualising the brain as an integrative network of functionally interacting brain regions offers a powerful framework for understanding large-scale neuronal communication. It provides a method for investigating how functional connectivity and information integration relates to human behaviour and how this organisation may be altered in neurodegenerative diseases (Bullmore & Sporns, 2009; Greicius et al., 2009).
@@
To understand how a specific brain region interacts with others, researchers most often analyse its resting-state activity and use simple pearson correlation of time-series data of a region with the time-series data of all other brain regions, they create a functional connectivity map (fcMap), which visually represents the strength of these connections (Biswal et al., 1997; Cordes et al., 2000). This is basically a matrix with value and we can understand it as a non relational data structure, in other terms, a graph.
@@
More and more work in cognitive neurosciences explore the link between graph theory and connectomes (functional connectivity matrices) (Farahani et al., 2019). By representing brain regions as nodes and their connections as edges, graph theory provides a powerful framework for analysing the structural and functional organisation of the brain. Notably, studies have begun to explore the link between structural properties of brain connectivity, as captured by connectomes, and the manifestation of neurological disorders such as Autism Spectrum Disorders (ASD) and Major Depressive Disorders (MDD). ASD, characterised by impairments in social communication and repetitive behaviours. MDD is characterised as a mood disorder marked by persistent sadness and loss of interest.
@@
These findings highlight the potential of connectome analysis to elucidate the neurological underpinnings of NDs and pave the way for the development of novel diagnostic and therapeutic strategies. Studying the link between the brain’s Functional connectivity signature and behavioural quality of patients through probing learned embeddings of neural networks trained on classification tasks could thus be a promising avenue to help disentangle the gap between its segregational characteristics and the emergence (Johnson, 2002; Eccles, 1994; Wang et al., 2015; Carroll & Parola, 2024) of higher level behavioural quality.
@@
However, if NDs result in alterations in brain functional and structural connections, as well as local and global connections (Seeley et al., 2009; Wang et al., 2015; Pasquini et al., 2015; Stam et al., 2007), traditional deep learning models such as CNN and LSTM are difficult to fit to the connectivity of the brain (Zhang et al., 2023). These long range dependencies, though, are well captured by the relational models defined previously in this thesis : Graph Neural Networks.
@@
Definition : Psychiatric diagnosis can be regarded as a graph classification task. Given an input graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with node feature matrix $X$, GNNs employ the message-passing paradigm to propagate and aggregate the representations of information along edges to generate a node representation $h_{v}$ for each node $v\in\mathcal{V}$ and then explore the modelled human brains using graph methods to extract abnormal brain networks, subnetworks, and local connections (Palop et al., 2006; Thomas et al., 2016).
@@
Similarly to (Zheng et al., 2023), a GNN can be formally defined through an aggregation function A and a combine function C such that $h_{v}^{(k)}$ is the node embedding of node $v$ at the $k$-th layer and $\mathcal{N}(v)$ is the set of neighbour nodes of $v$:
@@
In the context of connectomes, many studies have focused on the relationship between general intellectual ability and small-world characteristics in intrinsic functional networks for describing individual differences in general intelligence (van den Heuvel & Hulshoff Pol, 2010; van den Heuvel et al., 2009; Langer et al., 2012; Hilger et al., 2017). Better intellectual performance was associated with shorter characteristic path length, the nodal centrality of hub regions in the salience network, as well as the efficiency of functional integration between the frontal and parietal areas (Jung & Haier, 2007) In general, when connections between specialised brain regions are disrupted, even within localised areas, the result is often functional impairment. This impairment is linked to atypical integration of activity across distributed brain networks (Ffytche & Catani, 2005; Catani et al., 2005). Characterising this impairment through the use of GNN could be one application of our probing pipeline. So far, GNNs have achieved promising diagnostic accuracy on autism spectrum disorder (ASD) (Rakhimberdina et al., 2020), schizophrenia (Rakhimberdina & Murata, 2020), bipolar disorder (BD) (Yang et al., 2019) and MDD (Zheng et al., 2023). We’ll focus on ASD and MDD. But here as in other graph related fields, research has highlighted we were lacking Interpretability (Zheng et al., 2023).
For ASD, The contribution of rs-fMRI studies based on graph theory for autism exploration is important (Redcay et al., 2013; Rudie et al., 2013; Di Martino et al., 2014; Keown et al., 2017; Kazeminejad & Sotero, 2019). Studies have found increased short-range connections in ASD, particularly within sensory and association cortices. This local overconnectivity may contribute to the sensory sensitivities and restricted interests often seen in ASD. Conversely, long-range connections between distant brain regions tend to be reduced in ASD. This underconnectivity affects integration of information across brain networks. Based on this literature (Farahani et al., 2019) we know that the modularity, clustering coefficient, and local efficiency are relatively reduced in ASD (i.e., inefficiency of information transmission in a particular module) while global communication efficiency is increased (shorter average path lengths). As another example, (Redcay et al., 2013) observed an increase in betweenness centrality and local connections by analysing the prefrontal brain areas in adolescents with ASD.
@@
For patients with MDD, several studies have reported topological changes in human brain connectome, including a loss of the small-world network (Ye et al., 2015; Achard & Bullmore, 2007)] and a significant reorganisation of the community structure (Zhang et al., 2011; Leistedt et al., 2009; Lord et al., 2012). In general, MDD patients exhibit increased global and local clustering coefficients, indicating a higher degree of local interconnectedness and efficiency in information processing. Moreover, increased modularity in MDD patients indicated that there were relatively less inter-modular edges and more intra-modular edges, which may also be associated with the disruptions in emotion regulation by decreasing communications between the Default Mode Network (DMN) and the Cognitive Control Network (CCN) (Ye et al., 2015). We would thus expect that classifying FC matrices with regard to MDD should use more clustering coefficient, clusterization properties and modularity measures than random (like the presence of motifs like the number small clusters, squares or triangles).
G.1 Comparison between datasets